# Majorana fermion induced power-law scaling in the violations of the Wiedemann-Franz law

### v2

@@ Abstract
Violation of the Wiedemann-Franz (WF) law in a 2D topological insulator due to Majorana bound states (MBS) is studied via the Lorenz ratio in the single-particle picture. We study the scaling of the Lorenz ratio in the presence and absence of MBS with inelastic scattering modeled using a Buttiker voltage-temperature probe. We compare our results with that seen in a quantum dot junction in the Luttinger liquid picture operating in the topological Kondo regime. We explore the scaling of the Lorentz ratio in our setup when either inelastic scattering occurs with both phase and momentum relaxation or via phase relaxation alone. This scaling differs from that predicted by the Luttinger liquid picture for both uncoupled and coupled Majorana bound states and depends on the nature of inelastic scattering.
@@ I Introduction
Majorana fermions are a particular class of fermionic quasiparticles with the unique property of being their antiparticles. Majorana fermions have intrinsic topological protection from disorder. This property renders Majorana fermions immune to decoherence, making Majorana fermions an ideal candidate for quantum computation. Various proposals to generate and detect Majorana fermions exist within the field of condensed matter physics, such as semiconductor-superconductor heterostructures [1], normal metal-superconductor heterostructures, or topological insulator-superconductor heterostructures. There have been many attempts to detect Majorana fermions experimentally [2, 3, 4, 5], but there is still no irrefutable experimental evidence for the existence of Majorana fermions.
@@
There have been proposals to detect Majorana fermions using violations in WF law [6] by studying the scaling of the Lorenz ratio. Ref. [6] studies the violation of Wiedemann-Franz(WF) law in a quantum dot junction in the topological Kondo regime [7] hosting localized Majorana bound states (MBS). WF law states that the ratio of the electrical conductance to the thermal conductance is inversely proportional to the temperature. Majorana fermions are also known to break particle-hole symmetry (PHS) and induce violations of Wiedemann-Franz (WF) law [6]. In the presence of MBS, the Lorenz ratio has been shown to scale inversely with respect to the Luttinger parameter [6]. The setup in [6] is considered in the many body regime and shows that the Lorenz ratio shows power-law decay as a function of the Luttinger parameter in the setup.
@@
In this paper, we discuss the violation of WF law and the scaling of the Lorenz ratio in the presence of Majorana fermions using a Buttiker voltage-temperature probe [8, 9, 10, 11, 12], which induces inelastic scattering. We consider two kinds of inelastic scattering in our setup: phase and momentum relaxation [8, 13] and with only phase relaxation [9]. We study the scaling of the Lorenz ratio with respect to the strength of inelastic scattering and compare the scaling of the Lorenz ratio with the Luttinger parameter in the many-body setup considered in [6]. In our setup, we study three cases: (i) when MBS are absent, (ii) when MBS are present and uncoupled, and (iii) when MBS are coupled. When MBS are present the Lorenz ratio scales differently with inelastic scattering when both phase and momentum relaxation occur compared to when only phase relaxation occurs. When MBS are absent, WF law is not violated in presence of inelastic scattering. It is to be noted that there have been reports wherein WF law is shown to be violated due to chaotic transport, see Refs.[14, 15], suffice to say there is no chaotic transport in our work. The rest of the paper is organized as follows: In section II, we describe the motivation and importance of this work. Section III describes the scattering and transmission in our setup using Landauer-Buttiker scattering formalism [16, 17]. In section IV, we first calculate the thermoelectric coefficients like the Seebeck, Peltier, and thermal conductance in our setup using the Onsager relations. We then define the Lorenz ratio and introduce inelastic scattering via a Buttiker voltage-temperature probe [10]. In section V, we present an analysis of our results and compare the scaling of the Lorenz ratio with the strength of inelastic scattering and compare it with the scaling seen in Ref. [6]. We end with the conclusions in section VI.
@@ II Motivation
Ref [6] studies a quantum dot junction capable of hosting an even number of Majorana fermions to study the setup in the topological Kondo regime [7]. In special cases, the electrons in the quantum dot can couple to two-fold degenerate states. The regular Kondo effect [18] is a consequence of the coupling of mobile electrons in a confined region to spin-degenerate states. Further, an even number of Majorana fermions can couple non-locally, giving rise to two-fold degenerate states that can couple to electrons in the quantum dot [7]. It is known as the topological Kondo effect [7]. The setup is studied in the Luttinger liquid model, which uses many-body formalism and considers inelastic scattering due to electron-electron interaction. The electron-electron interaction is parameterized by the Luttinger parameter $(g)$, with $g=1$ corresponding to the absence of interaction, $g<1$ corresponding to attractive interaction, and $g>1$ corresponding to repulsive interaction. The authors of Ref. [6] show that the Lorenz ratio $(W)$ depends on the Luttinger parameter as $W(g)\approx\frac{2}{3g}$. When the setup in Ref. [6] is in the topological Kondo regime, the Majorana-induced boundary conditions and scattering via a splitting junction leads to the "splitting" of a charged particle into a transmitted particle of charge $\frac{2e}{3}$ and a backscattered hole of charge $\frac{e}{3}$. The coefficient $\frac{2}{3}$ is the unique and universal signature of Majorana bound states in a quantum dot junction operating in the topological regime according to Ref. [6].
@@
Our setup introduces inelastic scattering phenomenologically using Buttiker voltage probe [9, 8]. The Buttiker probe is an additional probe that induces inelastic scattering in the setup such that the total charge and heat current flowing into the Buttiker probe is zero. Unlike Luttinger liquid theory, the Buttiker voltage probe is based on single-particle scattering theory, making it more adaptable to setups studied using scattering theory. Further, using the Buttiker voltage probe, we can study inelastic scattering with phase and momentum relaxation [8] and inelastic scattering with only phase relaxation [9]. In our work, we look at both electric conductance and thermal conductance. Therefore, we modify the Buttiker voltage probe into a Buttiker voltage-temperature probe, as was also done recently in [10]. The comparison between the Luttinger parameter and inelastic scattering due to the Buttiker voltage-temperature probe [10] has not been studied before. We study the scaling of the Lorenz ratio (W) with respect to the coupling strength of the Buttiker voltage-temperature probe in our setup. A Buttiker voltage-temperature probe is added to the setup, such that the total charge and heat currents going into the probe vanish but induce inelastic scattering via both phase and momentum relaxation [8] or only phase relaxation [9]. Any electron/hole entering the Buttiker probe loses its phase memory and is reinjected with a completely different phase, leading to phase relaxation [9, 8]. For the case of both phase and momentum relaxation [8], the electrons injected into the probe from the setup are reinjected with equal probabilities of going toward the left or the right. Thus, the reinjection also causes the phase and the initial momentum to be lost. To preserve the momentum while inducing phase relaxation, one can use a pair of unidirectional probes such that when the electrons initially traveling to the left or the right are injected into the Buttiker probe, they are reinjected back with the same momentum, see Ref. [9]. We seek to probe MBS using two different kinds of Buttiker voltage-temperature probes, note the difference in these results, and compare our results with the results seen in Ref. [6], which is studied in the many body picture. This distinguishing behavior of MBS as a function of inelastic scattering in Buttiker voltage-temperature probe [10] is revealed through the Lorenz ratio. The signatures distinguish MBS’s existence and nature (coupled or individual).
@@ III.1 Hamiltonian
Figure 1: (a) The 3 terminal Majorana Aharonov-Bohm interferometer with a Buttiker voltage temperature probe, which entails both phase and momentum relaxation [8]. A 2D TI ring with helical edge modes flowing in the outer and inner edges. The ring is coupled to two regular TI leads via couplers (shown in orange). The leads connect the AB ring to the reservoirs. A third coupler connects the ring to an inelastic scatterer acting as a Buttiker voltage-temperature probe [10], such that the total charge and heat current flowing in the third terminal is zero. The ring is pierced by an Aharonov-Bohm flux. Ferromagnetic (shown in red) and superconducting (shown in cyan) correlations are induced in the top arm of the ring via the proximity effect. (b) The Majorana Aharonov-Bohm interferometer with a Buttiker voltage-temperature probe which entails only phase relaxation [9], the Buttiker voltage-temperature probe consists of two unidirectional probes attached to the bottom via a coupler such that the total charge and heat currents into each probe terminal is zero. The electrons injected into each probe terminal are reinjected back into the setup via the other probe, preserving the momentum. (c) The superconductor-topological insulator-ferromagnet (STIM) junction. MBS (shown as white ellipses) occurs at the interface of the ferromagnetic (shown in red) and superconducting (shown in cyan) layer. (d) Scattering of the edge modes at the left coupler. The solid and dashed lines represent the electron and hole edge modes. The black and white lines represent the spin-up and spin-down edge modes. The double-headed arrows represent the outer edge modes, and the single-headed arrows represent the inner edge modes, respectively.
@@ III.2 Transport in the system via edge modes
$S=\begin{pmatrix}-\frac{1}{3}I&&\frac{2}{3}I&&\frac{2}{3}I\\ &&&&\\ \frac{2}{3}I&&-\frac{1}{3}I&&\frac{2}{3}I\\ &&&\\ \frac{2}{3}I&&\frac{2}{3}I&&-\frac{1}{3}I\end{pmatrix},$ (4)
where $I$ is the $2\times 2$ identity matrix. This S-matrix is obtained by taking $\epsilon=\frac{4}{9}$ in Buttiker’s $3\times 3$ S-matrix, see Ref. [8]. The reason for using this S-matrix is that it corresponds exactly to a waveguide result, see Ref. [23]. While traversing the ring, the spin-up electrons and holes in the edge modes acquire a propagating phase [22] as follows:
@@
such that $l_{1}+l_{2}=l_{u}$, i.e. length of the upper arm of the ring. For the lower arm of the ring, left of the Buttiker voltage-temperature probe:
@@
for the lower arm of the ring, right of the Buttiker voltage-temperature probe:
@@
such that $l^{\prime}_{1}+l^{\prime}_{2}=l_{d}$, i.e., the length of the bottom arm of the ring. The total length of the ring $L$ is given by $L=l_{u}+l_{d}$. $k_{e}=(E+E_{f})/\hbar v_{F}$, and $k_{h}=(E-E_{f})/\hbar v_{F}$ are electron and hole wave vectors in the 2D TI. $\phi$ is the Aharonov-Bohm flux taken in units of the flux quantum $\phi_{0}=hc/e$. One can similarly find the phase acquired by the spin-down electrons and holes, which is the exact mirror image of spin-up electrons and holes. In the next section, we describe thermoelectric transport in multi-terminal mesoscopic systems and introduce inelastic scattering by adding a Buttiker voltage-temperature probe [8, 9, 10] to the Majorana ABI.
IV Inelastic scattering and Buttiker voltage-temperature probe
@@
Ls;kij= (Ls;kij;cVLs;kij;cTLs;kij;qVLs;kij;qT)= 1h∫-∞dE (δij- Tss;kkij(E, EF)+ Trs;lkij(E, EF)) ×ξ(E, EF)L0(E, EF),    with   r ≠s   and   l ≠k
with (1 (E - EF)/eTi(E-EF)/e (E-EF)2/e2Ti),
where $\delta_{ij}$ is the Dirac Delta function, $f(E,E_{F})$ being the Fermi function $(f(E,E_{F})=\frac{1}{1+e^{(E-E_{F})/k_{B}T_{i}}}$), with $\xi(E,E_{F})=\frac{-\partial f(E,E_{F})}{\partial E}$, $T_{1}$ and $T_{1}+\Delta T_{R}$ are the temperatures of the left and right reservoirs. In contrast, $T_{1}+\Delta T$ is the temperature of the Buttiker voltage-temperature probe, $k_{B}$ is the Boltzmann constant, $G_{0}=(e^{2}/\hbar)$ is the conductance quantum, $E$ is particle energy, $E_{F}$ is Fermi energy, and $h$ is Planck’s constant. The Onsager matrix elements describe the thermal and electrical response to the voltage and temperature bias. $L^{s;k}_{ij;cV}$, is the electrical conductance, $L^{s;k}_{ij;cT}$ is the electrical response to the temperature difference, $L^{s;k}_{ij;qV}$ is the thermal response to the voltage difference, and $L^{s;k}_{ij;qT}$ is the thermal response generated due to the temperature difference due to a spin $s$ electron or hole ($s\in\{\uparrow,\downarrow\}$) being transmitted from the $j^{th}$ terminal into the $i^{th}$ terminal as a spin $s$ electron or hole ($k\in\{e,h\}$). $\mathcal{T}_{ij}^{ss,kk}$ is the transmission probability for a particle of type $k\in\{e,h\}$ with spin $s\in\{\uparrow,\downarrow\}$ to transmit from terminal $j$ to terminal $i$ as the same particle type $k$ and same spin $s$. Similarly, $\mathcal{T}_{ij}^{sr,lk}$ is the transmission probability for a particle of type $k\in\{e,h\}$ with spin $s\in\{\uparrow,\downarrow\}$ to transmit from terminal $j$ to terminal $i$ as the particle type $l\neq k$ and spin $r\neq s$. MBS is a superposition of electrons and holes of the same spin; thus, the scattering due to MBS can cause electron-electron scattering with same spin or electron-hole scattering between particles of the opposite spin. Since the opposite spin edge modes are mirror images of each other and there is no spin-flip scattering in our system (see Fig. 1), we can write $\mathcal{T}^{\uparrow\uparrow;kk}_{ij}=\mathcal{T}^{\downarrow\downarrow;kk}_{ij}$, and $\mathcal{T}^{\downarrow\uparrow;lk}_{ij}=\mathcal{T}^{\uparrow\downarrow;kl}_{ij}$ and $\mathcal{T}^{\uparrow\downarrow}_{ij;kk}=\mathcal{T}^{\uparrow\downarrow}_{ij;kk}=0$. From Eq. (9), we can write:
@@
This section will introduce inelastic scattering in our system using a Buttiker voltage-temperature probe [10]. In the setup shown in Fig. 1, an additional lead is connected to the ABI via a coupler that connects the ABI to a Buttiker voltage-temperature probe such that the total charge current and the total heat current flowing into terminal 3 are zero. The S-matrices for the left and right couplers are described in Eq. (4), and they scatter electrons/holes elastically. The voltage-temperature probe induces inelastic scattering by setting the total charge and heat current passing through itself to zero. It ensures that no net current flows out of the setup into the probe (in section IV, analysis, we will look at the scaling of the Lorenz ratio with respect to the strength of inelastic scattering in the single-particle picture). We study two types of inelastic scattering in our setup, with only phase relaxation [9] and with both phase relaxation and momentum relaxation [8]. We compare and contrast our result with that obtained using Luttinger formalism [6].

### v4

@@ Abstract
Violation of the Wiedemann-Franz law in a 2D topological insulator due to Majorana bound states is studied via the Lorenz ratio in the single-particle picture. We study the scaling of the Lorenz ratio in the presence and absence of Majorana bound states with inelastic scattering modeled using a Buttiker voltage-temperature probe. We compare our results with that seen in a quantum dot junction in the Luttinger liquid picture operating in the topological Kondo regime. We explore the scaling of the Lorentz ratio in our setup when either phase and momentum relaxation or phase relaxation is present. This scaling differs from that predicted by the Luttinger liquid picture for both uncoupled and coupled Majorana cases..
@@ I Introduction
Majorana fermions are a particular class of fermions with the unique property of being their antiparticles. Majorana fermions have intrinsic topological protection from disorder. This property renders Majorana fermions immune to decoherence, making Majorana fermions an ideal candidate for quantum computation. These particles manifest as a zero-energy mode in the $p_{x}+ip_{y}$ wave superconductor as initially explained by Kitaev [1]. Moreover, their presence has been anticipated in 2D topological insulators interfaced with $s$-wave superconductors and separated by a ferromagnet [2, 3]. Numerous promising theoretical proposals have been suggested for the detection of Majorana bound states (MBS). By examining the conductance, distinctions between topological and trivial superconductors can be made, aiding in MBS detection [4, 5]. Hanbury-Brown and Twiss (HBT) non-local shot noise-like correlations [6], along with non-local conductance [7], have proven effective in distinguishing between topological and trivial superconductors, offering a reliable method for detecting MBS. Similarly, a Josephson junction modeled by superconductor-topological insulator-superconductor can exhibit a $4\pi$ periodic current-phase relationship instead of $2\pi$, attributed to MBS [8]. Further, in semiconductor-superconductor heterostructures [9], MBS are shown to occur. There have been numerous attempts to experimentally detect Majorana fermions [10, 11, 12, 13], but irrefutable evidence for their existence remains elusive. In [10], Majorana zero modes were initially reported based on the observation of a zero-bias conductance peak (ZBCP) quantized at $2e^{2}/h$ in a semiconductor nanowire coupled with an $s$-wave superconductor. However, this claim was later retracted. Similarly, Ref. [11] claimed the detection of Majorana zero modes in a hybrid semiconductor-superconductor interface by studying the Aharonov-Bohm effect and weak antilocalization, but this claim was also retracted. Other experimental works [12, 13] have also focused on ZBCPs as potential indicators of Majorana fermions, although these peaks alone may not definitively confirm the presence of Majorana bound states (MBS). In a recent study [14], a signature of crossed Andreev reflection was observed in a magnetic topological insulator that hosts the quantum anomalous Hall effect when in proximity to an $s$-wave superconductor such as Niobium. This signature of crossed Andreev reflection holds considerable significance for ongoing studies on Majorana bound states. There are some theoretical proposals on the connection between Majorana fermions and quantum metrology too [15].
@@
There have been proposals to detect Majorana fermions using violations in Wiedemann-Franz law [16] by studying the scaling of the Lorenz ratio. The Wiedemann-Franz (WF) law states that the ratio of effective thermal conductance to effective electrical conductance is proportional to the temperature of the system. In our case, the system is an Aharonov-Bohm ring composed of a topological insulator. We define this ratio as the Lorentz ratio. Ref. [16] studies the violation of Wiedemann-Franz law in a quantum dot junction in the topological Kondo regime [17] hosting localized Majorana bound states (MBS). WF law states that the ratio of the electrical conductance to the thermal conductance is inversely proportional to the temperature. The constant of proportionality, also known as the Lorenz ratio, is a constant for all conductors. Majorana fermions are also known to break particle-hole symmetry (PHS) and induce violations of Wiedemann-Franz (WF) law [16]. In the presence of MBS, the Lorenz ratio has been shown to scale inversely with respect to the Luttinger parameter [16]. The setup in [16] is considered in the many body regime and shows that the Lorenz ratio shows power-law decay as a function of the Luttinger parameter in the setup.
@@
In this paper, we discuss the violation of WF law and the scaling of the Lorenz ratio in the presence of Majorana fermions using a Buttiker voltage temperature (BVT) probe [18, 19, 20, 21, 22], which induces inelastic scattering. We consider two kinds of inelastic scattering in our setup: phase and momentum relaxation [18, 23] and with only phase relaxation [19]. We study the scaling of the Lorenz ratio with respect to the strength of inelastic scattering and compare the scaling of the Lorenz ratio with the Luttinger parameter in the many-body setup considered in [16]. In our setup, we study three cases: (i) when MBS are absent, (ii) when MBS are present and uncoupled, and (iii) when MBS are coupled. In a single-electron setup, the Lorenz ratio scales differently with inelastic scattering when both phase and momentum relaxation occur and when only phase relaxation occurs. For individual MBS, WF law is not violated without inelastic scattering.
The rest of the paper is organized as follows: In section II, we describe the motivation and importance of this work. Section III describes the scattering and transmission in our setup using Landauer-Buttiker scattering formalism [24, 25]. In section IV, we first calculate the thermoelectric coefficients like the Seebeck, Peltier, and thermal conductance in our setup using the Onsager relations. We then define the Lorenz ratio and introduce inelastic scattering via a BVT probe [20]. In section V, we present an analysis of our results and compare the scaling of the Lorenz ratio with the strength of inelastic scattering and compare it with the scaling seen in Ref. [16]. We end with the conclusions in section VI.
@@ II Motivation
The primary motivation of our work is indeed to detect Majorana-bound states through the observation of a power-law scaling in the violation of the Wiedemann-Franz law. However, our study also aims to explore whether this power-law scaling, previously observed in the Luttinger liquid regime, is also seen using a phenomenological model of BVT probe with edge mode transport in Landaeur-Buttiker scattering theory. This investigation is crucial because the many-body effects that are fundamental in the Luttinger liquid are phenomenologically incorporated into the scattering framework via the BVT probe. Our results demonstrate that while the nature of the violation of the Wiedemann-Franz law with respect to the inelastic scattering parameter in a single-particle picture remains similar between the two regimes, the power-law scaling differs. Notably, this work represents the only study where a violation of the Wiedemann-Franz law adheres to a specific scaling law within the Landauer-Buttiker scattering framework, offering a detailed comparison with the Luttinger liquid regime [16].
@@
Ref [16] studies a quantum dot junction capable of hosting an even number of Majorana fermions to study the setup in the topological Kondo regime [17]. In special cases, the electrons in the quantum dot can couple to two-fold degenerate states. The regular Kondo effect [26] is a consequence of the coupling of mobile electrons in a confined region to spin-degenerate states. Further, an even number of Majorana fermions can couple non-locally, giving rise to two-fold degenerate states that can couple to electrons in the quantum dot [17]. It is known as the topological Kondo effect [17]. The setup is studied in the Luttinger liquid model, which uses many-body formalism and considers inelastic scattering due to electron-electron interaction. The electron-electron interaction is parameterized by the Luttinger parameter $(g)$, with $g=1$ corresponding to the absence of interaction, $g<1$ corresponding to attractive interaction, and $g>1$ corresponding to repulsive interaction. The authors of Ref. [16] show that the Lorenz ratio $(W)$ depends on the Luttinger parameter as $W(g)=\frac{2}{3g}$. When the setup in Ref. [16] is in the topological Kondo regime, the Majorana-induced boundary conditions and scattering via a splitting junction leads to the "splitting" of a charged particle into a transmitted particle of charge $\frac{2e}{3}$ and a backscattered hole of charge $\frac{e}{3}$. The coefficient $\frac{2}{3}$ is the unique and universal signature of Majorana bound states in a quantum dot junction operating in the topological regime according to Ref. [16].
@@
Our setup introduces inelastic scattering phenomenologically using Buttiker voltage probe [19, 18]. The Buttiker probe is an additional probe that induces inelastic scattering in the setup such that the total charge and heat current flowing into the Buttiker probe is zero. Unlike Luttinger liquid theory, the Buttiker voltage probe is based on single-particle scattering theory, making it more adaptable to setups studied using scattering theory. Further, using the Buttiker voltage probe, we can study inelastic scattering with phase and momentum relaxation [18] and inelastic scattering with only phase relaxation [19]. In our work, we look at both electric conductance and thermal conductance. Therefore, we modify the Buttiker voltage probe into a BVT probe, as was also done recently in [20]. The comparison between the Luttinger parameter and inelastic scattering due to the BVT probe [20] has not been studied before. We study the scaling of the Lorenz ratio (W) with respect to the coupling strength of the BVT probe in our setup. A BVT probe is added to the setup, such that the total charge and heat currents going into the probe vanish but induce inelastic scattering via both phase and momentum relaxation [18] or only phase relaxation [19]. Any electron/hole entering the Buttiker probe loses its phase memory and is reinjected with a completely different phase, leading to phase relaxation [19, 18]. For the case of both phase and momentum relaxation [18], the electrons injected into the probe from the setup are reinjected with equal probabilities of going toward the left or the right. Thus, the reinjection also causes the phase and the initial momentum to be lost. To preserve the momentum while inducing phase relaxation, one can use a pair of unidirectional probes such that when the electrons initially traveling to the left or the right are injected into the Buttiker probe, they are reinjected back with the same momentum, see Ref. [19]. We seek to probe MBS using two different kinds of BVT probes, note the difference in these results, and compare our results with the results seen in Ref. [16], which is studied in the many body picture. Generally, in mesoscopic devices, noise and environmental effects [27] are modeled phenomenologically via the BVT probe in the setup with both phase and momentum relaxation and phase relaxation. This distinguishing behavior of MBS as a function of inelastic scattering in BVT probe [20] is revealed through the Lorenz ratio. The signatures distinguish MBS’s existence and nature (coupled or individual).
@@ III.1 Hamiltonian
Refer to captionJ1J2J1J2J1
Figure 1: (a) The 3 terminal Majorana Aharonov-Bohm interferometer with a BVT probe, which entails both phase and momentum relaxation [18]. A 2D topological insulator ring with helical edge modes flowing in the outer and inner edges. The ring is coupled to two regular topological insulator leads via couplers J1 and J2 (shown in orange). The leads connect the AB ring to the reservoirs. A third coupler connects the ring to an inelastic scatterer acting as a BVT probe [20], such that the total charge and heat current flowing in the third terminal is zero. The ring is pierced by an Aharonov-Bohm flux. Ferromagnetic (shown in red) and superconducting (shown in cyan) correlations are induced in the top arm of the ring via the proximity effect. (b) The Majorana Aharonov-Bohm interferometer with a BVT probe which entails only phase relaxation [19], the BVT probe consists of two unidirectional probes attached to the bottom via a coupler such that the total charge and heat currents into each probe terminal is zero. The electrons injected into each probe terminal are reinjected back into the setup via the other probe, preserving the momentum. (c) The superconductor-topological insulator-ferromagnet junction. Majorana bound states (shown as white ellipses) occurs at the interface of the ferromagnetic (shown in red) and superconducting (shown in cyan) layer. (d) Scattering of the edge modes at the left coupler J1. The solid and dashed lines represent the electron and hole edge modes. The black and white lines represent the spin-up and spin-down edge modes. The double-headed arrows represent the outer edge modes, and the single-headed arrows represent the inner edge modes, respectively.
@@ III.2 Transport in the system via edge modes
$S=\begin{pmatrix}-\left(\sqrt{1-2\alpha}\right)I&&\sqrt{\alpha}I&&\sqrt{\alpha}I\\ &&&&\\ \sqrt{\alpha}I&&\frac{1}{2}\left(\sqrt{1-2\alpha}-1\right)I&&\frac{1}{2}\left(\sqrt{1-2\alpha}-1\right)I\\ &&&\\ \sqrt{\alpha}I&&\frac{1}{2}\left(\sqrt{1-2\alpha}+1\right)I&&\frac{1}{2}\left(\sqrt{1-2\alpha}-1\right)I\end{pmatrix},$ (4)
where $I$ is the $2\times 2$ identity matrix. In our calculation, we take $\alpha=\frac{1}{2}$, however to analyze the plots of elastic scattering at couplers in Analysis in Sec. VI, we also consider $\alpha=\frac{4}{9}$ too in Eq. 4. While traversing the ring, the spin-up electrons and holes in the edge modes acquire a propagating phase [31] as follows:
@@
such that $l_{1}+l_{2}=l_{u}$, i.e. length of the upper arm of the ring. For the lower arm of the ring, left of the BVT probe:
@@
for the lower arm of the ring, right of the BVT probe:
@@
such that $l^{\prime}_{1}+l^{\prime}_{2}=l_{d}$, i.e., the length of the bottom arm of the ring. The total length of the ring $L$ is given by $L=l_{u}+l_{d}$. $k_{e}=(E+E_{f})/\hbar v_{F}$, and $k_{h}=(E-E_{f})/\hbar v_{F}$ are electron and hole wave vectors in the 2D TI. $\phi$ is the Aharonov-Bohm flux taken in units of the flux quantum $\phi_{0}=hc/e$. One can similarly find the phase acquired by the spin-down electrons and holes, which is the exact mirror image of spin-up electrons and holes. In the next section, we describe thermoelectric transport in multi-terminal mesoscopic systems and introduce inelastic scattering by adding a BVT probe [18, 19, 20] to the Majorana ABI.
IV Inelastic scattering and Buttiker voltage temperature probe
@@
[…]
where $\delta_{ij}$ is the Dirac Delta function, $f(E,E_{F})$ being the Fermi function $(f(E,E_{F})=\frac{1}{1+e^{(E-E_{F})/k_{B}T_{i}}}$), with $\xi(E,E_{F})=\frac{-\partial f(E,E_{F})}{\partial E}$, $T_{1}$ and $T_{1}+\Delta T_{R}$ are the temperatures of the left and right reservoirs. In contrast, $T_{1}+\Delta T$ is the temperature of the BVT probe, $k_{B}$ is the Boltzmann constant, $G_{0}=(e^{2}/\hbar)$ is the conductance quantum, $E$ is particle energy, $E_{F}$ is Fermi energy, and $h$ is Planck’s constant. The Onsager matrix elements describe the thermal and electrical response to the voltage and temperature bias. $L^{s;k}_{ij;cV}$, is the electrical conductance, $L^{s;k}_{ij;cT}$ is the electrical response to the temperature difference, $L^{s;k}_{ij;qV}$ is the thermal response to the voltage difference, and $L^{s;k}_{ij;qT}$ is the thermal response generated due to the temperature difference due to a spin $s$ electron or hole ($s\in\{\uparrow,\downarrow\}$) being transmitted from the $j^{th}$ terminal into the $i^{th}$ terminal as a spin $s$ electron or hole ($k\in\{e,h\}$). $\mathcal{T}_{ij}^{ss,kk}$ is the transmission probability for a particle of type $k\in\{e,h\}$ with spin $s\in\{\uparrow,\downarrow\}$ to transmit from terminal $j$ to terminal $i$ as the same particle type $k$ and same spin $s$. Similarly, $\mathcal{T}_{ij}^{sr,lk}$ is the transmission probability for a particle of type $k\in\{e,h\}$ with spin $s\in\{\uparrow,\downarrow\}$ to transmit from terminal $j$ to terminal $i$ as the particle type $l\neq k$ and spin $r\neq s$. MBS is a superposition of electrons and holes of the same spin; thus, the scattering due to MBS can cause electron-electron scattering with same spin or electron-hole scattering between particles of the opposite spin. Since the opposite spin edge modes are mirror images of each other and there is no spin-flip scattering in our system (see Fig. 1), we can write $\mathcal{T}^{\uparrow\uparrow;kk}_{ij}=\mathcal{T}^{\downarrow\downarrow;kk}_{ij}$, and $\mathcal{T}^{\downarrow\uparrow;lk}_{ij}=\mathcal{T}^{\uparrow\downarrow;kl}_{ij}$ and $\mathcal{T}^{\uparrow\downarrow}_{ij;kk}=\mathcal{T}^{\uparrow\downarrow}_{ij;kk}=0$. From Eq. (9), we can write:
@@
This section will introduce inelastic scattering in our system using a BVT probe [20]. In the setup shown in Fig. 1, an additional lead is connected to the ABI via a coupler that connects the ABI to a BVT probe such that the total charge current and the total heat current flowing into terminal 3 are zero. The S-matrices for the left and right couplers are described in Eq. (4), and they scatter electrons/holes elastically. The BVT probe induces inelastic scattering by setting the total charge and heat current passing through itself to zero. It ensures that no net current flows out of the setup into the probe (in section IV, analysis, we will look at the scaling of the Lorenz ratio with respect to the strength of inelastic scattering in the single-particle picture). We study two types of inelastic scattering in our setup, with only phase relaxation [19] and with both phase relaxation and momentum relaxation [18]. We compare and contrast our result with that obtained using Luttinger formalism [16].